# Polarization Diagnostics Applied to Coronal Mass Ejections and the Background Solar Wind

### v1

@@ Appendix A Definitions
  • $\varepsilon$ = elongation angle of line of sight (LOS) measured from X axis
  • $\alpha$ = polar angle of intersection of LOS with Y-Z plane (Observing Plane of Sky), measured counterclockwise from Z axis about X axis.
@@
  • $r_{pos}$ = distance from Sun center to intersection of LOS with Y-Z plane
@@
  • $d$ = distance of closest approach to Sun center of any LOS (impact factor; occurs at intersection of LOS and TS)
@@
  • $\tau$ = angle along LOS; $\tau=0$ is intersection of LOS with Thomson Sphere (TS), positive toward observer, negative away.
@@
  • $s$ = distance from TS to scattering point for a given LOS (positive toward observer, negative away)
@@ Appendix B Croissant Model
$\zeta$ is a spatial weighting ranging from zero to one: a two-dimensional cross-section is shown in Figure 26e, and a one-dimensional cross-section is plotted in Figure 26f. We set the total mass of the CME, and distribute that mass appropriately by $\zeta$. Furthermore, we apply a smooth function of height that distributes more of the mass at the CME apex compared to the footpoints. Without this function the CME legs dominate the brightness when FORWARD applies line-of-sight integrations. In principle, more realistic density distributions may be sought, although CME masses and density distributions are currently poorly constrained by observations.

### v2

@@ Appendix A Definitions
  • $\varepsilon$ = elongation angle of line of sight measured from X axis
  • $\alpha$ = polar angle of intersection of line of sight with Y-Z plane (Observing Plane of Sky), measured counterclockwise from Z axis about X axis.
@@
  • $r_{pos}$ = distance from Sun center to intersection of line of sight with Y-Z plane
@@
  • $d$ = distance of closest approach to Sun center of any line of sight (impact factor; occurs at intersection of line of sight and TS)
@@
  • $\tau$ = angle along line of sight; $\tau=0$ is intersection of line of sight with TS, positive toward observer, negative away.
@@
  • $s$ = distance from TS to scattering point for a given line of sight (positive toward observer, negative away)
@@ Appendix B Croissant Model
$\zeta$ is a spatial weighting ranging from zero to one: a two-dimensional cross-section is shown in Figure 26e, and a one-dimensional cross-section is plotted in Figure 26f. We set the total mass of the CME, and distribute that mass appropriately by $\zeta$. Furthermore, we apply a smooth function of height that distributes more of the mass at the CME apex compared to the footpoints. Without this function the CME legs dominate the brightness when FORWARD applies LOS integrations. In principle, more realistic density distributions may be sought, although CME masses and density distributions are currently poorly constrained by observations.